calibration(2f)technical.html[Stand09.F ebruar2014], year = 2011, and_{w}riting_{i}mages_{a}nd_{v}ideo.html?highlight = imwriteimwrite[Stand21.F ebruar2014], sensors/mt9m001c12stm/[Stand09.F ebruar2014],

¨ o ü ö ü ü date = January2011, volume =

3, number = 1, IGNOREmonth = jan, year =

2011, issn = 1941 − 6237, pages = 1 − −13, numpages =

13, REMurl = http : //dx.doi.org/10.4018/jaci.2011010101, doi = 10.4018/jaci.2011010101, acmid = 2433177, publisher = IGIGlobal, address = Hershey, P A, USA, keywords = BuiltEnvironmentNetworks(BENs), HumanP resence, NodeP lacement, RadiationP atterns, W irelessSensorNetworks(W SNs),

2012, author = IEEE, journal = IEEEStd802.11 − 2012(RevisionofIEEEStd802.11 − 2007), IGNOREtitle = StandardforInformationtechnology − −T elecommunicationsandinformationexchangebetweensystemsLocalandmetropolitanareanetworks − −Specificrequirements − P art11 : W irelessLANMediumAccessControl(MAC)andP hysicalLayer(P HY)Specifications, title = W irelessLANMediumAccessControl(MAC)andP hysicalLayer(P HY)Specifications, organization =

InstituteofElectricalandElectronicsEngineers, Inc., IGNOREmonth = 3, year = 2012, pages =

1 − 2793, publisher = IEEEComputerSociety, address = NewY ork, NY, USA, doi = 10.1109/IEEEST D.2012.6178212,

ž ä ä date = June1996, volume =

2, number =

2, IGNOREmonth = 6, year =

1996, issn = 1022 − 0038, pages = 143 − −154, numpages =

12, REMurl = http : //dx.doi.org/10.1007/BF 01225637, doi = 10.1007/BF 01225637, acmid = 234826, publisher = Springer − V erlagNewY ork, Inc., address =

Secaucus, NJ, USA, date = November2012, volume =

30, number =

4, IGNOREmonth = 11, year =

2012, issn = 1046 − 8188, IGNOREpages =

26 : 1 − −26 : 36, articleno = 26, numpages =

36, REMurl = http : //doi.acm.org/10.1145/2382438.2382445, doi = 10.1145/2382438.2382445, acmid = 2382445, publisher = ACM, address = NewY ork, NY, USA, keywords =

Opinion, blog, proximity, retrieval, sentiment, date = September2006, volume =

17, number =

7, IGNOREmonth = sep, year =

2006, issn = 0733 − 8716, pages = 1205 − −1211, numpages =

7, REMurl = http : //dx.doi.org/10.1109/49.778178, doi = 10.1109/49.778178, acmid = 2313013, publisher = IEEEP ress, address = P iscataway, NJ, USA, ö ü ü ö ä ö ö ü ö é ää 10_{1}2.pdf[] 58, title = 3DBloodF lowReconstructionfrom2DAngiograms, url = http : //dx.doi.org/10.1007/978 − 3 − 540 − 78640 − 5_{5}8, publisher = SpringerBerlinHeidelberg, author = P latzer, Esther − S.andDeinzer, F rankandP aulus, DietrichandDenzler, Joachim, pages = 288 − 292, language = English

é é Blanco_{P} hd_{T} hesis ü é ü ü ü ü \

ˆ {\} ˆ {\} \ \